# Sparse Representations of Dynamical Networks:

### v4

@@ 1. $\mathbf{b)}$
In contrast to the approach from [17], our technique places little emphasis on the sparsity pattern belonging to the realization of type (4a)-(4b), employed in the enforcement of sparsity patterns on the resulting SRTR pairs. Furthermore, as will be shown in the numerical example from Section 5, the proposed technique can produce SRTR pairs with distinctive network topologies and high degrees of sparsity, even when the realization employed by the procedure possesses a greatly reduced degree of sparsity. In spite of this, the arguments made in [17] remain relevant even in the considered case, which is to say that the discussed problem’s feasibility is highly dependent upon the network’s intrinsic flow of information. Thus, the aforementioned example will employ the procedure from [27] to obtain a controller with just such an information flow, imparted by a structured NRF representation.
Before moving on to the numerical example, we first discuss the computational aspects of the algorithms used to satisfy conditions $i)$ through $vi)$ from Theorem 4.25. Firstly, the orthogonal matrices $Q_{i}$ with $i\in 1:p$, which compress the individual rows of $A_{12}$, are straightforward to compute via the use of permutations and Householder transformations. Additionally, notice that conditions $i)$, $iv)$ and $v)$ only amount to a set of equality constraints in terms of linear combinations of the entries belonging to $K$, whereas condition $ii)$ must only be checked against the desired sparsity pattern.
@@
It has been shown via the identity from (80) in Appendix D that $B_{1}$ is invariant with respect to the class of all minimal realizations of a system showcasing the structure from (4a)-(4b). Therefore, the choice of minimal realization to be employed in Theorem 4.25 does not restrict the result’s applicability and means of inducing sparsity in $B_{1}$ will be presented in the numerical example from Section 5, by employing the synthesis procedures from [27], as previously indicated in Remark 4.28.
We conclude that the main difficulty in applying Theorem 4.25 rests with satisfying conditions $iii)$ and $vi)$. Notably, the former involves a series of equality constraints containing bilinear terms of the entries which belong to the matrix $K$, whereas the latter requires that we restrict the eigenvalues of an affine expression of $K$ to $\mathbb{S}$. Fortunately, these types of problems have been jointly tackled in Section IV of [27] and Algorithm 1 therein can easily be adapted to accommodate the conditions laid out in Theorem 4.25. While the procedures from [27] deal with continuous-time versions of condition $vi)$ from the aforementioned theorem, we point out that the discrete-time case can be handled just as elegantly, by employing Lemma 6 from [17] alongside the convexification techniques given in [27]. Alternatively, one may employ Lemma 8 from [17] to restrict the eigenvalues of $A_{22}+KA_{12}$ to $\mathbb{S}$ which, combined with condition $v)$ of Theorem 4.25, is sufficient to satisfy condition $vi)$ of the aforementioned theorem.
@@
The high computational performance of the indicated algorithm, as highlighted in Table 1 of [27] and discussed comparatively in Section V.B of [27], serves to justify the chosen approach and its distinctive formulation, given in Theorem 4.25. Furthermore, a code implementation for this type of numerical procedure is publicly available, and can be accessed via the link in the footnote from the first page of [27].
@@ 5 Numerical example
We now present an example of SRTR-based design for distributed control. To this end, we return to the network illustrated in Fig. 1 and modeled in Section 3.3. Recall that, although this network possesses a highly particular topology, the latter’s structure is completely obscured when inspecting the network via its TFM. The aim of this section is to illustrate how the framework developed in this paper can be used to obtain a controller, for the aforementioned network, which has the same ring topology and which can be implemented in a distributed manner, via a set of sparse and stabilizing control laws. The chosen example, inspired by the one showcased in Section V of [27], will be presented in the continuous-time setting. We point out that, in doing so, our aim is to provide further contrast with respect to the framework discussed in [35], which caters exclusively to discrete-time systems.
@@
where $\mathbf{Y}_{i}(s)$ and $\mathbf{U}_{i}(s)$ denote the Laplace transforms of the network’s scalar outputs and inputs, respectively, while the interconnection and input dynamics are given by the scalar TFMs $\mathbf{\Phi}_{\mathbf{G}}(s)$ and $\mathbf{\Gamma}_{\mathbf{G}}(s)$, respectively. Notice that these two scalar TFMs do not depend upon $i\in 1:p$, which indicates homogeneous dynamics among the network’s nodes (see [25] for a practical scenario where such a situation can occur).
@@
The observation pertaining to the state vector of the realization from (31) is crucial. By devising distributed control laws which stabilize the state dynamics of the system given in (31) in closed-loop configuration, this controller’s distributed implementation will guarantee that the states of all the $\mathbf{\Phi}_{\mathbf{G}}(s)$ and $\mathbf{\Gamma}_{\mathbf{G}}(s)$ subsystems forming the network will remain bounded when they are excited by bounded exogenous signals (as a consequence of the closed-loop state dynamics being made asymptotically stable, see Section 5.3 in [40] and Corollary III.9 from [24], along with its proof), and that these states (along with the state variables of the controller’s implementation) will tend to zero when evolving freely from any finite initial conditions.
In order to produce the control laws described by the identities from (28), we first employ the numerical procedures from [27] to obtain a controller whose NRF representation has the desired type of sparsity structure. This controller is given by a minimal state-space realization of order $n=12$ and type (4a)-(4b) denoted by
@@
Additionally, it is desirable to find low-order state-space realizations for each row of the TFM given in (22a) and, therefore, we choose $n_{i}=1,\ \forall i\in 1:6$. Finally, recall that we wish to obtain an SRTR-based control law in which each command signal is computed via the expression given in (28). For the controller produced by the procedures from [27], we obtain precisely these types of control laws when, in addition to the constraints stated in Theorem 4.25, we also impose that $A_{22}+KA_{12}$ be diagonal and that it have the same element on each position of its diagonal.
@@
(33) $\begin{array}[]{ll}\mathbf{W}_{local}(s)=\dfrac{-5.255s-55.9}{s+9.34}\ ,&\mathbf{V}_{local}(s)=\dfrac{-1.078s-94.28}{s+9.34}\ ,\\ \phantom{}\\ \mathbf{W}_{prev}(s)\hskip 0.7113pt=\dfrac{-15.84}{s+9.34}\ ,&\mathbf{V}_{prev}(s)\hskip 0.7113pt=\dfrac{15.84s-15.84}{s+9.34}\ .\\ \end{array}$
@@ 6 Conclusions
In this paper, we have introduced the SRTR representation for networked systems and we have presented a comprehensive discussion on the intrinsic connections between SRTR pairs, their NRF counterparts introduced in [24], the response TFMs from [35] with their associated implementation matrices in [13], and the consecrated coprime factorizations treated extensively in [31]. Akin to our previous results from [24], we have provided a unitary mathematical framework that is not restricted to either continuous- (see [10]) or discrete-time (see [35]) network models, while also highlighting the proposed paradigm’s rich potential for system analysis. Finally, we have extended the closed-loop internal stability guarantees from [24] to the current setting, thus paving the way for distributed SRTR-based implementations of control laws and validating our proposed framework’s capacity for controller design.
@@ Appendix A Rosenbrock-type realizations
We state here several results related to a type of realization which has been fully investigated in [21]. These remarkable representations are the generalizations of state-space [36], descriptor [30] and pencil [19, 20] realizations, being instrumental in constructing many of the proofs given in Appendix D.
@@
[21, pp. 111, Theorem 4.1] Let $\mathbf{G}(\lambda)$ be given by an irreducible realization of type (34). Then:
@@ Appendix B Non-symmetrical Algebraic Riccati Equations
We show here that the right stabilizing solutions of (19), as introduced in Definition 4.9, can be readily expressed and reliably computed by employing a particular invariant subspace (see Section 1.3 of [9]) of the matrix
@@
Let $\begin{bmatrix}V_{1}^{\top}&V_{2}^{\top}\end{bmatrix}^{\top}$ have $w$ columns and full column rank, for some integer $w\in 1:n$, with $V_{1}$ having $p$ rows and $V_{2}$ having $n-p$ rows. By point 2 of Theorem 1.3.2 in [9], we have that if
@@
[…]
@@
Proposition B.3 (Corollary 4.2 in [11]).
The CTNARE from (19) has a right stabilizing solution $K$ if and only if $A^{+}$ has a disconjugate $\mathbb{S}$-invariant subspace of dimension $p$. In such cases, if we denote any basis for such a subspace as $\begin{bmatrix}V_{1}^{\top}&V_{2}^{\top}\end{bmatrix}^{\top}$, then $V_{1}$ is invertible and $K=V_{2}V_{1}^{-1}$ is a right stabilizing solution of (19).
@@ Appendix C Auxiliary Results
and notice that, due to Assumption 4.1, $\begin{bmatrix}S_{1}^{\top}(\lambda)&S_{2}^{\top}(\lambda)\end{bmatrix}^{\top}$ has full row rank $\forall\lambda\in\mathbb{C}$ and $\forall K\in\mathbb{R}^{(n-p)\times p}$. Then, $S_{2}(\lambda)$ must share this property, which makes the realization from (15) controllable. Since this realization is also observable, we get that it is minimal and that all the eigenvalues of $A_{22}+KA_{12}$ are the poles of its TFM. To be stable, these eigenvalues must be located in $\mathbb{S}$, thus proving the statement’s necessity.
The following result is a classical state-space derivation, originating from [18] and generalized in [15] (see also Proposition 1 in [15]). It proves crucial in establishing the connections between stable SRTR pairs and LCFs over $\mathbb{S}$, in Sections 4.2 and 4.3.
@@ 3. $\mathbf{c)}$
(40) $\begin{bmatrix}\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{cc|cc}A_{x}-\lambda I_{n-p}&B_{x}A_{12}&A_{x}B_{x}-B_{x}A_{11}+B_{x}A_{12}K&B_{x}B_{1}\\ O&A_{22}+KA_{12}-\lambda I_{p}&KA_{12}K+A_{22}K-KA_{11}-A_{21}&KB_{1}+B_{2}\\ \hline\cr C_{x}&O&C_{x}B_{x}&O\\ \end{array}\right]$
@@
Let $\mathbf{G}_{1}\in\mathbb{R}(\lambda)^{p\times p}$ and $\mathbf{G}_{2}\in\mathbb{R}(\lambda)^{p\times m}$ such that we can express the following realization of order $n-p>0$, given explicitly by
@@
(43) $\overline{\mathbf{G}}(\lambda):=\big{(}I_{p}-\lambda^{-1}\mathbf{G}_{1}(\lambda)\big{)}^{-1}\big{(}\lambda^{-1}\mathbf{G}_{2}(\lambda)\big{)}=\left[\scriptsize\begin{array}[]{cc|c}{D}_{1}-\lambda I_{p}&{C}&{D}_{2}\\ {B}_{1}&{A}-\lambda I_{n-p}&{B}_{2}\\ \hline\cr I_{p}&O&O\end{array}\right].$
@@
(60) $=$ $\footnotesize\left[\begin{array}[]{cc|c}-\lambda I_{p}&{C}&{D}_{1}\\ O&{A}-\lambda I_{n-p}&{B}_{1}\\ \hline\cr-I_{p}&O&I_{p}\end{array}\right]^{-1}\footnotesize\left[\begin{array}[]{cc|c}-\lambda I_{p}&{C}&{D}_{1}\\ O&{A}-\lambda I_{n-p}&{B}_{1}\\ \hline\cr I_{p}&O&O\end{array}\right]$
@@
(72) $\overline{\mathbf{G}}(\lambda)=\left[\scriptsize\begin{array}[]{cccc|c}{D}_{1}-\lambda I_{p}&{C}&O&O&{D}_{2}\\ {B}_{1}&{A}-\lambda I_{n-p}&O&O&{B}_{2}\\ O&O&-\lambda I_{p}&{C}&{D}_{2}\\ O&O&O&{A}-\lambda I_{n-p}&{B}_{2}\\ \hline\cr I_{p}&O&O&O&O\end{array}\right].$
@@ Appendix D Proofs of the main results
(73) $\overline{\mathbf{G}}(\lambda):=\big{(}I_{p}-\lambda^{-1}\mathbf{W}(\lambda)\big{)}^{-1}\big{(}\lambda^{-1}\mathbf{V}(\lambda)\big{)}=\left[\tiny\begin{array}[]{cc|c}A_{11}-A_{12}K-\lambda I_{p}&A_{12}&B_{1}\\ A_{K}&A_{22}+KA_{12}-\lambda I_{n-p}&KB_{1}+B_{2}\\ \hline\cr I_{p}&O&O\end{array}\right],$
@@
In addition to the equality from (74), notice that $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ are expressed via state-space realizations, which means that they are guaranteed to be proper TFMs. Moreover, from the expression of $\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$ given in (15), we get that the latter TFM’s McMillan degree is at most $n-p$. thus, by Definition 3.8, any $\big{(}\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big{)}$ obtained via (15) is indeed an SRTR pair of the system from the result’s statement.
To prove point $\mathbf{b)}$, we first show that minimality is necessary for (15) to produce, given the plant’s realization, every SRTR pair associated with the latter realization and having a McMillan degree of at most $n-p$. Let us assume that this is not the case and consider the system described by the realization of type (4a)-(4b)
@@
To show the sufficiency of starting from a system given by minimal realization (4a)-(4b) of order $n$, having $\mathbf{G}(\lambda)$ as its TFM, we will first consider an arbitrary SRTR pair $\big{(}\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big{)}$ associated with the aforementioned system. Thus, by Definition 3.8, we can always obtain a realization of order $n-p$ for the following TFM
(78) $\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{c|cc}\overline{A}-\lambda I_{n-p}&\overline{B}_{1}&\overline{B}_{2}\\ \hline\cr\vspace{-2mm}\hfil\\ \overline{C}&\overline{D}_{1}&\overline{D}_{2}\end{array}\right],\vspace{-1mm}$
@@
(80) $\mathbf{G}(\lambda)=\left[\scriptsize\begin{array}[]{cc|c}A_{11}-A_{12}K-\lambda I_{p}&A_{12}L^{-1}&B_{1}\\ L(KA_{11}-KA_{12}K+A_{21}-A_{22}K)&L(A_{22}+KA_{12})L^{-1}-\lambda I_{n-p}&L(KB_{1}+B_{2})\\ \hline\cr I_{p}&O&O\end{array}\right].$
@@
(81) $\scriptsize\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{c|cc}L(A_{22}+KA_{12})L^{-1}-\lambda I_{n-p}&L(KA_{11}-KA_{12}K+A_{21}-A_{22}K)&L(KB_{1}+B_{2})\\ \hline\cr A_{12}L^{-1}&A_{11}-A_{12}K&B_{1}\\ \end{array}\right]\normalsize.$
@@
(82) $\begin{bmatrix}\lambda I_{p}-\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}-\lambda\widetilde{E}&\widetilde{B}-\lambda\widetilde{F}\\ \hline\cr\widetilde{C}&\widetilde{D}\end{array}\right]:=\\ :=\left[\footnotesize\begin{array}[]{cc|cc}A_{22}+KA_{12}-\lambda I_{n-p}&O&KA_{12}K+A_{22}K-KA_{11}-A_{21}&KB_{1}+B_{2}\\ O&I_{p}&-\lambda I_{p}&O\\ \hline\cr A_{12}&I_{p}&-A_{11}+A_{12}K&B_{1}\\ \end{array}\right].\normalsize$
@@
(83) $\mathcal{S}(\lambda):=\scriptsize\begin{bmatrix}A_{22}+KA_{12}-\lambda I_{n-p}&O&KA_{12}K+A_{22}K-KA_{11}-A_{21}&KB_{1}+B_{2}\\ O&I_{p}&-\lambda I_{p}&O\\ A_{12}&I_{p}&-A_{11}+A_{12}K&B_{1}\\ \end{bmatrix}.\normalsize\vspace{-1mm}$
@@
$\footnotesize\begin{bmatrix}\widetilde{A}-\lambda\widetilde{E}\\ \widetilde{C}\end{bmatrix}=\footnotesize\begin{bmatrix}A_{22}+KA_{12}-\lambda I_{n-p}&O\\ O&I_{p}\\ A_{12}&I_{p}\end{bmatrix}\normalsize=\footnotesize\begin{bmatrix}I_{n-p}&-K&K\\ O&I_{p}&O\\ O&O&I_{p}\end{bmatrix}\footnotesize\begin{bmatrix}A_{22}-\lambda I_{n-p}&O\\ O&I_{p}\\ A_{12}&I_{p}\end{bmatrix}\vspace{-1mm}$

### v7

@@ 1. $\mathbf{b)}$
In contrast to the approach from [Realiz], our technique places little emphasis on the sparsity pattern belonging to the realization of type (4a)-(4b), employed in the enforcement of sparsity patterns on the resulting SRTR pairs. Furthermore, as will be shown in the numerical example from Section 5, the proposed technique can produce SRTR pairs with distinctive network topologies and high degrees of sparsity, even when the realization employed by the procedure possesses a greatly reduced degree of sparsity. In spite of this, the arguments made in [Realiz] remain relevant even in the considered case, which is to say that the discussed problem’s feasibility is highly dependent upon the network’s intrinsic flow of information. Thus, the aforementioned example will employ the procedure from [aug_sparse] to obtain a controller with just such an information flow, imparted by a structured NRF representation.
Before moving on to the numerical example, we first discuss the computational aspects of the algorithms used to satisfy conditions $i)$ through $vi)$ from Theorem 4.25. Firstly, the orthogonal matrices $Q_{i}$ with $i\in 1:p$, which compress the individual rows of $A_{12}$, are straightforward to compute via the use of permutations and Householder transformations. Additionally, notice that conditions $i)$, $iv)$ and $v)$ only amount to a set of equality constraints in terms of linear combinations of the entries belonging to $K$, whereas condition ii)ii) must only be checked against the desired sparsity pattern.
@@
It has been shown via the identity from (80) in Appendix D that $B_{1}$ is invariant with respect to the class of all minimal realizations of a system showcasing the structure from (4a)-(4b). Therefore, the choice of minimal realization to be employed in Theorem 4.25 does not restrict the result’s applicability and means of inducing sparsity in $B_{1}$ will be presented in the numerical example from Section 5, by employing the synthesis procedures from [aug_sparse], as previously indicated in Remark 4.28.
We conclude that the main difficulty in applying Theorem 4.25 rests with satisfying conditions $iii)$ and $vi)$. Notably, the former involves a series of equality constraints containing bilinear terms of the entries which belong to the matrix $K$, whereas the latter requires that we restrict the eigenvalues of an affine expression of $K$ to $\mathbb{S}$. Fortunately, these types of problems have been jointly tackled in Section IV of [aug_sparse] and Algorithm 1 therein can easily be adapted to accommodate the conditions laid out in Theorem 4.25. While the procedures from [aug_sparse] deal with continuous-time versions of condition $vi)$ from the aforementioned theorem, we point out that the discrete-time case can be handled just as elegantly, by employing Lemma 6 from [Realiz] alongside the convexification techniques given in [aug_sparse]. Alternatively, one may employ Lemma 8 from [Realiz] to restrict the eigenvalues of $A_{22}+KA_{12}$ to $\mathbb{S}$ which, combined with condition $v)$ of Theorem 4.25, is sufficient to satisfy condition $vi)$ of the aforementioned theorem.
@@
The high computational performance of the indicated algorithm, as highlighted in Table 1 of [aug_sparse] and discussed comparatively in Section V.B of [aug_sparse], serves to justify the chosen approach and its distinctive formulation, given in Theorem 4.25. Furthermore, a code implementation for this type of numerical procedure is publicly available, and can be accessed via the link in the footnote from the first page of [aug_sparse].
@@ 5 Numerical example
We now present an example of SRTR-based design for distributed control. To this end, we return to the network illustrated in Fig. 1 and modeled in Section 3.3. Recall that, although this network possesses a highly particular topology, the latter’s structure is completely obscured when inspecting the network via its TFM. The aim of this section is to illustrate how the framework developed in this paper can be used to obtain a controller, for the aforementioned network, which has the same ring topology and which can be implemented in a distributed manner, via a set of sparse and stabilizing control laws. The chosen example, inspired by the one showcased in Section V of [aug_sparse], will be presented in the continuous-time setting. We point out that, in doing so, our aim is to provide further contrast with respect to the framework discussed in [SLS], which caters exclusively to discrete-time systems.
@@
where $\mathbf{Y}_{i}(s)$ and $\mathbf{U}_{i}(s)$ denote the Laplace transforms of the network’s scalar outputs and inputs, respectively, while the interconnection and input dynamics are given by the scalar TFMs $\mathbf{\Phi}_{\mathbf{G}}(s)$ and $\mathbf{\Gamma}_{\mathbf{G}}(s)$, respectively. Notice that these two scalar TFMs do not depend upon $i\in 1:p$, which indicates homogeneous dynamics among the network’s nodes (see [plutonizare] for a practical scenario where such a situation can occur).
@@
The observation pertaining to the state vector of the realization from (31) is crucial. By devising distributed control laws which stabilize the state dynamics of the system given in (31) in closed-loop configuration, this controller’s distributed implementation will guarantee that the states of all the $\mathbf{\Phi}_{\mathbf{G}}(s)$ and $\mathbf{\Gamma}_{\mathbf{G}}(s)$ subsystems forming the network will remain bounded when they are excited by bounded exogenous signals (as a consequence of the closed-loop state dynamics being made asymptotically stable, see Section 5.3 in [zhou] and Corollary III.9 from [NRF], along with its proof), and that these states (along with the state variables of the controller’s implementation) will tend to zero when evolving freely from any finite initial conditions.
In order to produce the control laws described by the identities from (28), we first employ the numerical procedures from [aug_sparse] to obtain a controller whose NRF representation has the desired type of sparsity structure. This controller is given by a minimal state-space realization of order $n=12$ and type (4a)-(4b) denoted by
@@
Additionally, it is desirable to find low-order state-space realizations for each row of the TFM given in (22a) and, therefore, we choose $n_{i}=1,\ \forall i\in 1:6$. Finally, recall that we wish to obtain an SRTR-based control law in which each command signal is computed via the expression given in (28). For the controller produced by the procedures from [aug_sparse], we obtain precisely these types of control laws when, in addition to the constraints stated in Theorem 4.25, we also impose that $A_{22}+KA_{12}$ be diagonal and that it have the same element on each position of its diagonal.
@@
[…]
@@ 6 Conclusions
In this paper, we have introduced the SRTR representation for networked systems and we have presented a comprehensive discussion on the intrinsic connections between SRTR pairs, their NRF counterparts introduced in [NRF], the response TFMs from [SLS] with their associated implementation matrices in [Separ], and the consecrated coprime factorizations treated extensively in [V]. Akin to our previous results from [NRF], we have provided a unitary mathematical framework that is not restricted to either continuous- (see [Explicit]) or discrete-time (see [SLS]) network models, while also highlighting the proposed paradigm’s rich potential for system analysis. Finally, we have extended the closed-loop internal stability guarantees from [NRF] to the current setting, thus paving the way for distributed SRTR-based implementations of control laws and validating our proposed framework’s capacity for controller design.
@@ Appendix A Rosenbrock-type realizations
We state here several results related to a type of realization which has been fully investigated in [Rose70]. These remarkable representations are the generalizations of state-space [Wonham], descriptor [Verg81] and pencil [CF, SIMAX] realizations, being instrumental in constructing many of the proofs given in Appendix D.
@@
[Rose70, pp. 111, Theorem 4.1] Let $\mathbf{G}(\lambda)$ be given by an irreducible realization of type (34). Then:
@@ Appendix B Non-symmetrical Algebraic Riccati Equations
We show here that the right stabilizing solutions of (19), as introduced in Definition 4.9, can be readily expressed and reliably computed by employing a particular invariant subspace (see Section 1.3 of [genricc]) of the matrix
@@
Let $\begin{bmatrix}V_{1}^{\top}&V_{2}^{\top}\end{bmatrix}^{\top}$ have $w$ columns and full column rank, for some integer $w\in 1:n$, with $V_{1}$ having $p$ rows and $V_{2}$ having npn-p rows. By point 2 of Theorem 1.3.2 in [genricc], we have that if
@@
When $S$ has all its eigenvalues located in $\mathbb{S}$, the subspace being spanned by the basis from (37) is called $\mathbb{S}$-invariant.
@@
Proposition B.3 (Corollary 4.2 in [EJC]).
The CTNARE from (19) has a right stabilizing solution $K$ if and only if $A^{+}$ has a disconjugate $\mathbb{S}$-invariant subspace of dimension pp. In such cases, if we denote any basis for such a subspace as $\begin{bmatrix}V_{1}^{\top}&V_{2}^{\top}\end{bmatrix}^{\top}$, then $V_{1}$ is invertible and $K=V_{2}V_{1}^{-1}$ is a right stabilizing solution of (19).
@@ Appendix C Auxiliary Results
and notice that, due to Assumption 4.1, $\begin{bmatrix}S_{1}^{\top}(\lambda)&\hskip-5.69054ptS_{2}^{\top}(\lambda)\end{bmatrix}^{\top}$ has full row rank $\forall\lambda\in\mathbb{C}$ and $\forall K\in\mathbb{R}^{(n-p)\times p}$. Then, $S_{2}(\lambda)$ must share this property, which makes the realization from (15) controllable. Since this realization is also observable, we get that it is minimal and that all the eigenvalues of $A_{22}+KA_{12}$ are the poles of its TFM. To be stable, these eigenvalues must be located in $\mathbb{S}$, thus proving the statement’s necessity.
The following result is a classical state-space derivation, originating from [Nett] and generalized in [Lucic] (see also Proposition 1 in [Lucic]). It proves crucial in establishing the connections between stable SRTR pairs and LCFs over $\mathbb{S}$, in Sections 4.2 and 4.3.
@@ 3. $\mathbf{c)}$
(40) $\begin{bmatrix}\mathbf{M}(\lambda)&\hskip-5.69054pt\mathbf{N}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{cc|cc}\hskip-2.84526ptA_{x}-\lambda I_{n-p}&B_{x}A_{12}&A_{x}B_{x}-B_{x}A_{11}+B_{x}A_{12}K&B_{x}B_{1}\\ O&\hskip-5.69054ptA_{22}+KA_{12}-\lambda I_{p}&\hskip-2.84526ptKA_{12}K+A_{22}K-KA_{11}-A_{21}&\hskip-2.84526ptKB_{1}+B_{2}\\ \hline\cr C_{x}&O&C_{x}B_{x}&O\\ \end{array}\right]$
@@
Let $\mathbf{G}_{1}\in\mathbb{R}(\lambda)^{p\times p}$ and $\mathbf{G}_{2}\in\mathbb{R}(\lambda)^{p\times m}$ such that we can express the following realization of order np>0n-p>0, given explicitly by
@@
(43) $\overline{\mathbf{G}}(\lambda):=\big(I_{p}-\lambda^{-1}\mathbf{G}_{1}(\lambda)\big)^{-1}\big(\lambda^{-1}\mathbf{G}_{2}(\lambda)\big)=\left[\scriptsize\begin{array}[]{cc|c}\hskip-2.84526pt{D}_{1}-\lambda I_{p}&{C}&{D}_{2}\\ {B}_{1}&\hskip-2.84526pt{A}-\lambda I_{n-p}\hskip-2.84526pt&{B}_{2}\\ \hline\cr I_{p}&O&O\end{array}\right].$
@@
(60) $=$ $\footnotesize\left[\begin{array}[]{cc|c}-\lambda I_{p}&{C}&{D}_{1}\\ O&{A}-\lambda I_{n-p}&{B}_{1}\\ \hline\cr-I_{p}&O&I_{p}\end{array}\right]^{-1}\hskip-2.84526pt\footnotesize\left[\begin{array}[]{cc|c}-\lambda I_{p}&{C}&{D}_{1}\\ O&{A}-\lambda I_{n-p}&{B}_{1}\\ \hline\cr I_{p}&O&O\end{array}\right]$
@@
(72) $\overline{\mathbf{G}}(\lambda)=\hskip-1.42262pt\left[\scriptsize\begin{array}[]{cccc|c}\hskip-2.84526pt{D}_{1}-\lambda I_{p}&{C}&O&O&{D}_{2}\\ {B}_{1}&{A}-\lambda I_{n-p}&O&O&{B}_{2}\\ O&O&-\lambda I_{p}&{C}&{D}_{2}\\ O&O&O&{A}-\lambda I_{n-p}&{B}_{2}\\ \hline\cr I_{p}&O&O&O&O\end{array}\right].$
@@ Appendix D Proofs of the main results
(73) $\overline{\mathbf{G}}(\lambda):=\big(I_{p}-\lambda^{-1}\mathbf{W}(\lambda)\big)^{-1}\big(\lambda^{-1}\mathbf{V}(\lambda)\big)\hskip-1.42262pt=\hskip-1.42262pt\left[\tiny\begin{array}[]{cc|c}\hskip-2.84526ptA_{11}-A_{12}K-\lambda I_{p}&A_{12}&B_{1}\\ A_{K}&\hskip-2.84526ptA_{22}+KA_{12}-\lambda I_{n-p}\hskip-2.84526pt&\hskip-2.84526ptKB_{1}+B_{2}\hskip-2.84526pt\\ \hline\cr I_{p}&O&O\end{array}\right],$
@@
In addition to the equality from (74), notice that $\mathbf{W}(\lambda)$ and $\mathbf{V}(\lambda)$ are expressed via state-space realizations, which means that they are guaranteed to be proper TFMs. Moreover, from the expression of $\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}$ given in (15), we get that the latter TFM’s McMillan degree is at most npn-p. thus, by Definition 3.8, any $\big(\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big)$ obtained via (15) is indeed an SRTR pair of the system from the result’s statement.
To prove point $\mathbf{b)}$, we first show that minimality is necessary for (15) to produce, given the plant’s realization, every SRTR pair associated with the latter realization and having a McMillan degree of at most npn-p. Let us assume that this is not the case and consider the system described by the realization of type (4a)-(4b)
@@
To show the sufficiency of starting from a system given by minimal realization (4a)-(4b) of order $n$, having $\mathbf{G}(\lambda)$ as its TFM, we will first consider an arbitrary SRTR pair $\big(\mathbf{W}(\lambda),\mathbf{V}(\lambda)\big)$ associated with the aforementioned system. Thus, by Definition 3.8, we can always obtain a realization of order npn-p for the following TFM
(78) $\begin{bmatrix}\mathbf{W}(\lambda)&\mathbf{V}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{c|cc}\overline{A}-\lambda I_{n-p}&\overline{B}_{1}&\overline{B}_{2}\\ \hline\cr\vskip-5.69054pt\\ \overline{C}&\overline{D}_{1}&\overline{D}_{2}\end{array}\right],\vskip-2.84526pt$
@@
(80) $\mathbf{G}(\lambda)=\left[\scriptsize\begin{array}[]{cc|c}A_{11}-A_{12}K-\lambda I_{p}&A_{12}L^{-1}&B_{1}\\ \hskip-2.84526ptL(KA_{11}-KA_{12}K+A_{21}-A_{22}K)&L(A_{22}+KA_{12})L^{-1}-\lambda I_{n-p}&L(KB_{1}+B_{2})\\ \hline\cr I_{p}&O&O\end{array}\right].$
@@
(81) $\scriptsize\begin{bmatrix}\mathbf{W}(\lambda)&\hskip-5.69054pt\mathbf{V}(\lambda)\end{bmatrix}=\left[\scriptsize\begin{array}[]{c|cc}\hskip-2.84526ptL(A_{22}+KA_{12})L^{-1}-\lambda I_{n-p}&L(KA_{11}-KA_{12}K+A_{21}-A_{22}K)&\hskip-5.69054ptL(KB_{1}+B_{2})\\ \hline\cr A_{12}L^{-1}&A_{11}-A_{12}K&B_{1}\\ \end{array}\right]\normalsize\hskip-2.84526pt.$
@@
(82) $\begin{bmatrix}\lambda I_{p}-\mathbf{W}(\lambda)&\hskip-5.69054pt\mathbf{V}(\lambda)\end{bmatrix}=\left[\footnotesize\begin{array}[]{c|c}\widetilde{A}-\lambda\widetilde{E}&\widetilde{B}-\lambda\widetilde{F}\\ \hline\cr\widetilde{C}&\widetilde{D}\end{array}\right]:=\\ :=\left[\footnotesize\begin{array}[]{cc|cc}\hskip-5.69054ptA_{22}+KA_{12}-\lambda I_{n-p}&O&\hskip-2.84526ptKA_{12}K+A_{22}K-KA_{11}-A_{21}&\hskip-5.69054ptKB_{1}+B_{2}\hskip-2.84526pt\\ O&I_{p}&-\lambda I_{p}&O\\ \hline\cr A_{12}&I_{p}&-A_{11}+A_{12}K&B_{1}\\ \end{array}\right]\hskip-2.84526pt.\normalsize$
@@
[…]
@@
$\footnotesize\begin{bmatrix}\widetilde{A}-\lambda\widetilde{E}\\ \widetilde{C}\end{bmatrix}=\footnotesize\begin{bmatrix}A_{22}+KA_{12}-\lambda I_{n-p}&O\\ O&I_{p}\\ A_{12}&I_{p}\end{bmatrix}\normalsize=\footnotesize\begin{bmatrix}I_{n-p}&-K&K\\ O&I_{p}&O\\ O&O&I_{p}\end{bmatrix}\footnotesize\begin{bmatrix}A_{22}-\lambda I_{n-p}&O\\ O&I_{p}\\ A_{12}&I_{p}\end{bmatrix}\vskip-2.84526pt$